the cost of three note books and four pencils is $8.50.the cost of five notebooks and eight pencils is $14.50. determine the cost of one notebook and the cost of one pencil

Answer by checkley79(3341)

3N+4P=8.50 MULTIPLY BY -2 & ADD.
5N+8P=14.50
-6N-8P=-17.00
---------------------
-N=-2.50
N=$2.50 ANS. FOR THE PRICE OF 1 NOTEBOOK
3*2.50+4P=8.50
7.50+4P=8.50
4P=8.50-7.50
4P=1.00
P=1.00/4
P=.$25 ANS. FOR THE COST OF 1 PENCIL.
PROOF:
5*2.50+4*.25=8.50
7.50+1.00=8.50
8.50=8.50